OpenStudy (anonymous):

Find the slope of line AB for the given points describe the slant of the segment 1.A(1,1) B(5,4)

OpenStudy (anonymous):

Do you know the formula to find slope?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

The slope is \[\frac{(y_2 - y_1)}{(x_2 - x_1)}\] Where y_2 is the y-coordinate of one of the points, and x_2 is the x-coordinate of that same point, and the same with y_1 and x_1. You have to make sure they are in order or it'll come out wrong So \[\frac{(4 - 1)}{(5 - 1)} = \frac{3}{4}\] So there's your slope

OpenStudy (anonymous):

ty